# Learning Visual Generative Priors without Text

### v2

@@ 1 Introduction
Such a philosophy inspires us to study image-to-image (I2I) generation, where models can learn from in-the-wild images in a self-supervised manner. We first develop a pure vision-based framework, termed as Lumos, to generate images conditioned on image features extracted by a pre-trained vision encoder (e.g., DINO). Then, we fine-tune Lumos to various downstream visual synthesis tasks (e.g., T2I). As shown in LABEL:fig:intro_a, thanks to I2I prior, our T2I model shows less reliance on high-quality data, exhibiting 0.5 less degradation in CLIP score compared with w/o I2I prior. Moreover, as illustrated in LABEL:fig:intro_b, with the data of I2I scaling up from 10M to 200M, our model achieves better performance, confirming the feasibility and the scalability of learning I2I priors. We hypothesize that the performance gap arises because the I2I priors enhance texture modeling, thereby simplifying the T2I generative process. Another interesting question lies in the influence of image encoders on I2I model. For a T2I model, text-image alignment is crucial. Therefore, a straightforward approach would be to use a text-image-aligned CLIP in the I2I model, which may enhance the T2I model’s performance compared to DINO. However, as shown in LABEL:fig:intro_c, although CLIP shows greater advantages in the early stages of learning I2I priors, DINO, as a late bloomer, achieves better COCO-30K FID than CLIP in the final steps. It demonstrates the feasibility of the pure vision-based training framework. Based on these analyses, we find that, as an upstream task of T2I, our pure-vision I2I model serves as a more foundational visual prior.
@@ 4.1 Image-to-Image Generation
Implementation Details & Datasets. Since our image-to-image generation does not require labels, we can easily filter totaling 190 million images via some criteria of image quality from existing open-source datasets (i.e., LAION-5B [44], COYO-700M [5], SAM [26], JourneyDB [33], and ImageNet-1K [14]). More details can refer to Supplementary Material.
We use DINO-B as the vision encoder of the image-to-image model and DiT-XL-2 [34] as the diffusion backbone. We evaluate the image-to-image model via Fréchet Inception Distance (FID) [22] on MSCOCO [28] and ImageNet-1K [14].
@@
Comparison with SOTA I2I model. We compare our model with the open source state-of-the-art I2I model, i.e., RCG [27], which is pre-trained on ImageNet-1K [14]. For the RCG model, we use the open-source version, namely pre-trained DiT-XL/2 (the same as our diffusion model structure) conditioned on Moco-v3 ViT-B [20]. As shown in Table 1, benefiting from the larger image dataset, Lumos outperforms RCG in terms of FID on ImageNet-1K and MSCOCO, indicating that our generated images are more consistent with the distribution of real images. Moreover, we compare the performances of different types of vision encoder for the I2I task in Figure 9 (a).
@@ 4.2 Downstream Visual Generative Tasks
Subsequently, a pre-trained I2I model as visual generative prior is used to transfer to various downstream tasks, including Text-to-Image Generation, Novel View Synthesis, and Image-to-Video Generation.
@@ 4.2.1 Text-to-Image Generation
We extensively evaluate the performance of Lumos and compare it to the results of state-of-the-art T2I models.
Implementation Details & Datasets. We construct a dataset of 30 million text-image pairs, all of which are sourced from easily accessible open datasets. This includes 10M and 5M images selected from LAION-5B and COYO-700M respectively. Additionally, 10M from SAM [26], 4M from JourneyDB [33], and 1M from Imagenet-1K [14] are selected. Following DALLE-3 [2], we incorporate raw captions and detailed captions generated by InternVL [11] into training. Besides FID, we evaluate T2I models on GenEval [18] that focus on compositionality. We use the T5 [38] large language model (specifically 4.3B Flan-T5-XXL) as the text encoder of our T2I model. More details are in Supplementary Material.
Quantitative Comparison for FID score. In Table 2, we report the evaluation results on commonly used zero-shot MS-COCO dataset in terms of FID score, and compare Lumos with existing SOTA T2I models. It is worth mentioning that we evaluate the open-sourced model of PixArt-$\alpha$ [7] that is trained on all the datasets proposed in thier paper. Thanks to a pre-trained image-to-image model as a more foundational visual prior, our T2I model achieves on-par or better performance than existing T2I models using only 1/10 text-image pairs for fine-tuning. Due to that our T2I model use the long caption to train, we employ a multi-modal large language model to generate long captions for evaluated dataset of MSCOCO. With the help of long captions, our text-to-image model surpasses all current text-to-image models in terms of the FID metric.
Quantitative Comparison for Image-text Alignment Assessment. Beyond the image fidelity evaluation, we evaluate the alignment between the generated images and text condition using GenEval [18]. GenEval is a comprehensive benchmark, designed for evaluating the compositional text-to-image generation capability. It evaluates the generation ability of the text-to-image model from multiple perspectives, including the number of objects, color, attributes, and position. As shown in the comparison results of Table. 3, the performance of our model is on par with the existing models with less than one billion parameters.
Table 2: Comparison to the recent text-to-image models on COCO FID-30k. ‘T&I Pairs’ refers to the number of the training text-image pairs. refers to the result of “PixArt-$\alpha$-256”. indicates evaluating with the long captions. Both ’+’ and ’-’ in the table denote the unknown internal dataset size and training step.
@@
Table 3: Text-image alignment comparison with the state-of-the-art text-to-image models on GenEval. Our model is decently trained on publicly available datasets without using any self-collected data. We highlight the best and second best.
@@
SDXL [35] 2.6B 0.98 0.74 0.39 0.85 0.23 0.15 0.55
SD3-8B  [17] 8.0B 0.99 0.94 0.72 0.89 0.60 0.33 0.74
FLUX-dev [3] 12.0B 0.99 0.81 0.79 0.74 0.47 0.20 0.67
FLUX-schnell [3] 12.0B 0.99 0.92 0.73 0.78 0.54 0.28 0.71
SDv1.5 [42] 0.9B 0.97 0.38 0.35 0.76 0.06 0.04 0.43
SDv2.1 [42] 0.9B 0.98 0.51 0.44 0.85 0.17 0.07 0.50
PixArt-$\alpha$ [7] 0.6B 0.98 0.50 0.44 0.80 0.07 0.08 0.48
PixArt-$\Sigma$ [8] 0.6B 0.98 0.59 0.50 0.80 0.15 0.10 0.52
SD3-1B [17] 1.0B 0.97 0.72 0.52 0.78 0.34 0.16 0.58
Lumos-T2I 0.8B 0.99 0.64 0.52 0.84 0.30 0.15 0.57
@@ 4.2.2 Novel View Synthesis
We further validate the advantages of the image-to-image model in 3D tasks through the novel view synthesis task.
Implementation Details & Datasets. We finetune our pre-trained image-to-image models using a subset (750k) of the released Objaverse [13] dataset.Following Zero123 [29], we randomly sample 32 camera extrinsic matrices $\mathcal{M}_{\rceil}$ which are oriented towards the center of the object, followed by rendering 32 views using a ray tracing engine.
The novel view synthesis models are evaluated on Google Scanned Objects (GSO) [16], which is a dataset of high-quality scanned household objects. We evaluate the models extensively with three metrics covering different aspects of image similarity: PSNR, SSIM, and LPIPS. More details are in Supplementary Material.
I2I $v.s.$ T2I Prior on NVS Task. To validate the effectiveness of I2I prior to the novel view synthesis task, we compare with training from scratch, I2I prior, and T2I prior on the novel view synthesis task. For fair comparison, we use the SOTA method, i.e., PixArt-$\alpha$ [7] model, as T2I prior, whose structure is similar to our model. As shown in Figure 5, thank to I2I prior, the performance of our method is better than no prior and T2I prior. It demonstrates the superiority of I2I prior over T2I prior on novel view synthesis task.
Comparison with SOTA NVS models. We compare our novel view synthesis model with the existing state-of-the-art models on the GSO dataset. Since the SyncDreamer [30] can only generate new perspective photos with a 30 elevation, we uniformly render photos at 16 angles around the object at two elevations of 0 and 30 for testing. The comparison results of Table 4 show that benefiting from the I2I prior, our novel view synthesis model outperforms the current state-of-the-art models.
Table 4: Comparison with SOTA models on novel view synthesis task. ‘Sync.’ indicates the SyncDreamer [30] model, which can only generate novel views at elevation 30.
@@
Qualitative Results. Figure 6 illustrates a qualitative comparison of novel view generation results for a GSO test object. As can be seen, the novel view images we generated are more notably consistent and realistic.
Figure 6: Comparison qualitative results between Lumos with the SOTA models. The object in the leftmost column represents the reference camera viewpoint, and the novel view images generated by our model exhibit significant advantages in realism and consistency.
@@ 4.2.3 Image-to-Video Generation
Following the experimental setup in Stable Video Diffusion [4], we fine-tune our image-to-image model for the image-to-video (I2V) generation task. The image-to-video task takes a still-input image as a conditioning input to generate subsequent continuous video frames.
Implementation Details & Datasets. We use the common open-source video datasets Webvid-10M [1] and a subset of Openvid-1M [31], which contains approximately 0.4M videos. We report the Fréche Video Distance (FVD) [48], the Kernel Video Distance (KVD) [48], and the Perceptual Input Conformity (PIC) [52] on UCF-101 [45] and MSR-VTT [53]. More details are in Supplementary Material.
I2I $v.s.$ T2I Prior on I2V Task. We compare I2I and T2I prior for the image-to-video generation task through zero-shot evaluating on UCF-101 [45] and MSR-VTT [53] datasets. For the T2I prior, the experiment setting follows the ones of the NVS experiments. As shown in Table 5, T2I and I2I prior exhibit a notable initialization for I2V task. Meanwhile, the comparative results indicate that the I2I prior holds distinct advantages in this task.
@@ 4.3 Ablation Study
Additionally, we conduct ablation studies to evaluate the effectiveness of different model components for image-to-image prior on text-to-image task.
@@
Comparison to Different Priors. We compare our I2I prior with class-to-image prior and uncondition prior, and also compare it with the image-to-image prior from RCG. To ensure fair experiments, all frameworks are trained on ImageNet-1K [14] and the same image-text dataset. As shown in Figure 7 (a), the results demonstrate that fine-tuning with I2I prior has a faster convergence speed than other two priors on T2I task. Moreover, we compare our I2I framework with the existing I2I framework, i.e., RCG [27], that uses adaLN-Zero block. The results in Figure 7 (b) indicate that our framework provides better prior knowledge.
@@
Comparison between Different Types of Condition Features for I2I prior. To explore which types of visual features (global semantic or fine-grained local information) are more suitable for the I2I framework, we set up three comparative experiments, i.e., a single global token ([CLS] token), local patch tokens, and all vision tokens (single global token + local patch tokens). As shown in Figure 8 (a) and (b), sufficient fine-grained local information significantly accelerates the convergence speed during the pre-training process. However, it is interesting that fine-grained local features lead to a high dependence of the pre-trained model on the conditions, which is not conducive to the rapid transfer of the pre-trained model to downstream tasks.
@@
Comparison between Different Vision Encoders. We set up experiments to probe the impact of different types of vision encoders (e.g., unimodal self-supervised $v.s.$ multimodal alignment) on our I2I frameworks. We adopt three types of vision encoders: CLIP, DINO, and MoCoV3. The comparison results in Figure 9 (a) demonstrate that the self-supervised pre-trained vision encoder (i.e., DINO and MoCoV3) has a significantly faster convergence speed in the image-to-image task. We then validate the transfer performance on the downstream task that relies most on text-image alignment (text-to-image generation). The comparison results in Figure 9 (b) show that the image-text alignment models have significant advantages during the initialization phase of transfer learning. However, this advantage quickly diminishes during the fine-tuning process. Thus, I2I generation framework has a robustness for different vision encoder types.
@@
Effectiveness of Different Text Encoders for T2I task. We fairly compare the performance of T5 and CLIP text encoder in text-to-image task based on different I2I models, as shown in Figure 10 (a) and (b). In the I2I pre-training framework with unimodal vision encoders (e.g., DINO), T5 outperformed the CLIP text encoder in both convergence speed and performance in terms of FID. For multimodal vision encoders (e.g., CLIP-L), the CLIP text encoder shows greater advantages in the early stages of learning I2I priors. However, T5, as a late bloomer, achieves better FID than CLIP in the final steps.
@@
Effectiveness of Class-Free Guidance for I2I prior. We ablate condition dropout probability during the training process and the class-free guidance factor in inference phrase in Figure 11 (a) and (b). We investigate that the condition dropout probability during training exhibits similar properties in image-to-image tasks as in text-to-image tasks. However, for the class-free guidance factor, we observe that image generation tasks have a broader applicability and a significant robustness advantage.

### v3

@@ 1 Introduction
Such a philosophy inspires us to study image-to-image (I2I) generation, where models can learn from in-the-wild images in a self-supervised manner. We first develop a pure vision-based framework, termed as Lumos, to generate images conditioned on image features extracted by a pre-trained vision encoder (e.g., DINO). Then, we fine-tune Lumos to various downstream visual synthesis tasks (e.g., T2I). As shown in LABEL:fig:intro_a, thanks to I2I prior, our T2I model shows less reliance on high-quality data, exhibiting 0.5 less degradation in CLIP score compared with w/o I2I prior. Moreover, as illustrated in LABEL:fig:intro_b, with the data of I2I scaling up from 10M to 200M, our model achieves better performance, confirming the feasibility and the scalability of learning I2I priors. We hypothesize that the performance gap arises because the I2I priors enhance texture modeling, thereby simplifying the T2I generative process. Another interesting question lies in the influence of image encoders on I2I model. For a T2I model, text-image alignment is crucial. Therefore, a straightforward approach would be to use a text-image-aligned CLIP in the I2I model, which may enhance the T2I model’s performance compared to DINO. However, as shown in LABEL:fig:intro_c, although CLIP shows greater advantages in the early stages of learning I2I priors, DINO, as a late bloomer, achieves better COCO-30K FID than CLIP in the final steps. It demonstrates the feasibility of the pure vision-based training framework. Based on these, we find that, as an upstream task of T2I, our pure-vision I2I model serves as a more foundational visual prior.
@@ 4.1 Image-to-Image Generation
Implementation Details & Datasets. Since our image-to-image generation does not require labels, we can easily filter totaling 190 million images via some criteria of image quality from existing open-source datasets (i.e., LAION-5B [46], COYO-700M [5], SAM [27], JourneyDB [35], and ImageNet-1K [14]). More details can refer to Supplementary Material. We use DINO-B as the vision encoder of the image-to-image model and DiT-XL-2 [36] as the diffusion backbone. We evaluate the image-to-image model via Fréchet Inception Distance (FID) [22] on MSCOCO [30] and ImageNet-1K [14].
@@
Comparison with SOTA I2I model. We compare our model with the open-source state-of-the-art I2I model, i.e., RCG [29], which is pre-trained on ImageNet-1K [14]. For the RCG model, we use the open-source version, namely pre-trained DiT-XL/2 (the same as our diffusion model structure) conditioned on Moco-v3 ViT-B [20]. As shown in Table 1, Lumos significantly outperforms RCG in terms of FID on ImageNet-1K and MSCOCO, demonstrating the excellent scalability of Image-to-Image generation on large-scale datasets.
@@ 4.2 Downstream Visual Generative Tasks
Subsequently, Lumos-I2I is used as visual generative prior and transferred to various downstream tasks, including Text-to-Image Generation, Novel View Synthesis, and Image-to-Video Generation.
@@ 4.2.1 Text-to-Image Generation
We first transfer Lumos-I2I to the text-to-image generation task and compare Lumos-T2I with the state-of-the-art T2I models in terms of the quality of the generated images and alignment assessment.
Table 2: Comparison to the recent text-to-image models on COCO FID-30k. ‘ T&I Pairs ’ refers to the number of the training text-image pairs. refers to the result of ‘ PixArt-$\alpha$-256 ’. indicates evaluating with the long captions. Both + and - in the table denote the unknown internal dataset size and training step.
@@
Implementation Details & Datasets. We construct a dataset of 30 million text-image pairs, all of which are sourced from easily accessible open datasets. This includes 10M and 5M images selected from LAION-5B and COYO-700M, respectively. Additionally, 10M from SAM [27], 4M from JourneyDB [35], and 1M from Imagenet-1K [14] are selected. Following DALLE-3 [2], we incorporate raw captions and detailed captions generated by InternVL [11] into training. Besides FID, we evaluate T2I models on GenEval [18] and DPG-Bench [23]. We use the T5 [40] (specifically 4.3B Flan-T5-XXL) as the text encoder of our T2I model. More details are in Supplementary Material.
Quantitative Comparison for FID score. In Table 2, we report the evaluation results on the widely used MS-COCO dataset in terms of FID score and compare Lumos with existing SOTA T2I models. It is worth mentioning that we evaluate the open-sourced model of PixArt-$\alpha$ [7] that is trained on all the datasets proposed in their paper. Thanks to the visual generative prior from the pre-trained image-to-image model, Lumos-T2I model achieves on-par or better performance than existing T2I models, requiring only a minimal amount of training data and steps. Moreover, Lumos-T2I surpasses all current text-to-image models in terms of the FID metric with the help of long captions. The experimental results demonstrate the significant benefits of the I2I prior for T2I models.
Table 3: Text-image alignment comparison with the state-of-the-art text-to-image models on GenEval [18] and DPG-Bench [23]. Our model is decently trained on publicly available datasets without using any self-collected data. We highlight the best and second best.
@@
[…]
Single Two Attribution
LUMINA-Next [59] 2.0B 0.92 0.46 0.48 0.70 0.13 0.09 0.46 82.8 88.7 86.4 80.5 81.8 74.6
SDXL [37] 2.6B 0.98 0.74 0.39 0.85 0.23 0.15 0.55 83.3 82.4 80.9 86.8 80.4 74.7
Playground v2.5 [28] 2.6B 0.98 0.77 0.52 0.84 0.17 0.11 0.56 83.1 82.6 81.2 84.1 83.5 75.5
SD3-8B [17] 8.0B 0.99 0.94 0.72 0.89 0.60 0.33 0.74 - - - - - -
FLUX-dev [3] 12.0B 0.99 0.81 0.79 0.74 0.47 0.20 0.67 82.1 89.5 88.7 91.1 89.4 84.0
FLUX-schnell [3]
12.0B 0.99 0.92 0.73 0.78 0.54 0.28 0.71 91.2 91.3 89.7 86.5 87.0 84.8
SDv1.5 [44] 0.9B 0.97 0.38 0.35 0.76 0.06 0.04 0.43 74.6 74.2 75.4 73.5 67.8 63.2
SDv2.1 [44] 0.9B 0.98 0.51 0.44 0.85 0.17 0.07 0.50 77.7 78.1 74.9 80.7 80.7 68.1
PixArt-$\alpha$ [7] 0.6B 0.98 0.50 0.44 0.80 0.07 0.08 0.48 81.7 80.1 80.4 81.7 76.5 71.6
PixArt-$\Sigma$ [8] 0.6B 0.98 0.59 0.50 0.80 0.15 0.10 0.52 87.5 87.1 86.5 84.0 86.1 79.5
SD3-1B [17] 1.0B 0.97 0.72 0.52 0.78 0.34 0.16 0.58 - - - - - -
Lumos-T2I 0.8B 0.99 0.64 0.52 0.84 0.30 0.15 0.57 87.4 87.4 86.5 87.2 88.1 79.9
Quantitative Comparison for Image-text Alignment Assessment. Beyond the image fidelity evaluation, we evaluate the alignment between the generated images and text prompts through GenEval [18] and DPG-Bench [23]. GenEval [18] is designed for evaluating T2I models from multiple perspectives, including the number of objects, color, attributes, and position. DPG-Bench [23] is a more challenging benchmark, which evaluates the generating performance of the dense prompts and the relation between the entities. Despite using only a small amount of training data and steps, the ability of Lumos-T2I’s text-image alignment is on par with or exceeds existing models with a similar parameter scale, as shown in the comparison results of Table. 3. And on the more challenging DPG-Bench [23], Lumos-T2I’s performance is still outstanding. This further proves the feasibility of I2I pre-training followed by T2I alignment training framework in the text-to-image task.
@@ 4.2.2 Novel View Synthesis
We further validate the advantages of the I2I prior on 3D tasks through the novel view synthesis task.
Implementation Details & Datasets. We transfer Lumos-I2I to novel view synthesis through a subset (750k) of the released Objaverse [13] dataset. Following Zero123 [31], we randomly sample 32 camera extrinsic matrices $\mathcal{M}_{\rceil}$ which are oriented towards the center of the object, followed by rendering 32 views using a ray tracing engine. Lumos-NVS is evaluated on Google Scanned Objects (GSO) [16], which is a dataset of high-quality scanned household objects. We evaluate the models extensively with three metrics covering different aspects of image similarity: PSNR, SSIM, and LPIPS. More details are in Supplementary Material.
I2I $v.s.$ T2I Prior on NVS Task. To validate the effectiveness of I2I prior on the novel view synthesis task, we compare with training from scratch, I2I prior, and T2I prior on the novel view synthesis task. For fair comparison, we use the SOTA method, i.e., PixArt-$\alpha$ [7] model, as T2I prior, whose structure is similar to our model. As shown in Figure 4, the model transferred from I2I prior is better than from no prior and T2I prior, which demonstrates the superiority of I2I prior over T2I prior.
Comparison with SOTA NVS models. We compare Lumos-NVS with the existing state-of-the-art models in  Table 4. We uniformly render photos at 16 angles around the object at two elevations of 0 and 30 for testing. The comparison results show that benefiting from the I2I prior, Lumos-NVS outperforms the current state-of-the-art models.
Table 4: Comparison with SOTA models on novel view synthesis task.
@@
Qualitative Results. Figure 6 illustrates a qualitative comparison of novel view generation results for a GSO test object. As can be seen, the novel view images Lumos-NVS generated are more notably consistent and realistic.
Figure 6: Comparison of qualitative results between Lumos-NVS with the SOTA models.
@@ 4.2.3 Image-to-Video Generation
Following the experimental setup in Stable Video Diffusion [4], we transfer Lumos-I2I to Lumos-I2V for the image-to-video (I2V) generation task. The image-to-video task takes a still-input image as a conditioning input to generate subsequent continuous video frames.
Implementation Details & Datasets. We use the common open-source video datasets Webvid-10M [1] and a subset of Openvid-1M [33], which contains approximately 0.4M videos. For evaluating the performance of I2V models, we report the Fréche Video Distance (FVD) [50], the Kernel Video Distance (KVD) [50], and the Perceptual Input Conformity (PIC) [54] on UCF-101 [47] and MSR-VTT [55]. More details are in Supplementary Material.
I2I $v.s.$ T2I Prior on I2V Task. We compare I2I and T2I prior for the image-to-video generation task through zero-shot evaluating on UCF-101 [47] and MSR-VTT [55] datasets. For the T2I prior, the experiment setting follows the ones of the NVS experiments. As shown in Table 5, T2I and I2I prior exhibit a notable initialization for the I2V task. Meanwhile, the comparative results indicate that the I2I prior holds distinct advantages in this task.
@@ 4.3 Ablation Study
Finally, we conduct a comprehensive series of experiments to thoroughly analyze and evaluate the effectiveness of various generative priors and model components.
@@
Ablating different generative priors. We compare Lumos-I2I prior with the class-to-image prior and uncondition prior. We also conduct a comparison with I2I prior from RCG [29], which uses adaLN-Zero block for injecting conditions. To ensure fair experiments, all frameworks are pre-trained on ImageNet-1K [14] and fine-tuned on the same image-text dataset. As shown in Figure 7 (a) and (b), Lumos-I2I provides better prior knowledge for downstream the T2I generation task.
@@
Ablating different features for I2I generative prior. To explore which types of visual features (global semantic or fine-grained local information) are more suitable for the I2I framework, we set up three comparative experiments, i.e., a single global token ([CLS] token), local patch tokens, and all vision tokens (single global token + local patch tokens). As shown in Figure 8 (a) and (b), sufficient fine-grained local information significantly accelerates the convergence speed during the I2I training process. However, it is interesting that fine-grained local features lead to a high dependence of the pre-trained model on the conditions, which is not conducive to the rapid transfer of the pre-trained model to downstream tasks.
@@
Ablating vision encoders for I2I generative prior. We set up experiments to probe the impact of different types of vision encoders (e.g., unimodal self-supervised $v.s.$ multimodal alignment) on our I2I frameworks. We adopt three types of vision encoders: CLIP, DINO, and MoCoV3. The comparison results in Figure 9 (a) demonstrate that the self-supervised pre-trained vision encoder (i.e., DINO and MoCoV3) has a significantly faster convergence speed in the image-to-image task. We then validate the transfer performance on the downstream task that relies most on text-image alignment (text-to-image generation). The comparison results in Figure 9 (b) show that the image-text alignment models have significant advantages during the initialization phase of transfer learning. However, this advantage quickly diminishes during the fine-tuning process. Thus, I2I generation framework has a robustness for different vision encoder types.
@@
Ablating different text Encoders for T2I task. We compare the performance of T5 and CLIP text encoder on different I2I generative priors, as shown in Figure 10 (a) and (b). For the I2I generative priors with unimodal vision encoders (e.g., DINO), T5 outperformed the CLIP text encoder in both convergence speed and performance in terms of FID. For priors with multimodal vision encoders (e.g., CLIP-L), the corresponding CLIP text encoder shows greater advantages in the early transferring stage. However, T5 finally achieves better performance.
Ablating class-free guidance for I2I model. We ablate condition dropout probability during the training process and the class-free guidance factor in the inference phrase in Figure 11 (a) and (b). We investigate that the condition dropout probability during training exhibits similar properties in image-to-image tasks as in text-to-image tasks. However, for the class-free guidance factor, we observe that image generation tasks have a broader applicability and a significant robustness advantage.